# The 1979 Iranian Revolution and the Lost Decade of Science: A Counterfactual Scientometric Analysis

### v1

@@ 3.6 Cross-country synthesis on a log scale (Figure 10)
Before 1979, Iran’s slope is competitive—at times steeper than South Korea’s—and its level is comparable to Taiwan’s, indicating genuine pre-1979 momentum from a nontrivial base. Immediately after 1979, Iran’s curve flattens and dips, signalling a collapse in compounding; by contrast, South Korea, Taiwan, and Singapore enter multi-decadal high-slope regimes during the 1980s–1990s. This divergence in rates explains why Iran’s strong 2000s recovery does not translate into convergence with the Tigers.
Panel B (middle): Iran vs. established/regional scientific systems (Netherlands, Israel, Greece).
The Netherlands and Israel trace mature, high-level trajectories with moderately rising slopes through the period—consistent with large, stable research systems that compound steadily rather than explosively. Greece shows a long run-up from a lower base, accelerating after European integration. Relative to this group, Iran’s post-1979 dip is distinctive: the level falls below Greece in the 1980s and only re-approaches the regional band after 2000. The key contrast is regime stability: the mature systems preserve low implied doubling times throughout the 1980s–1990s, whereas Iran shifts into a high-doubling-time regime and only partially reverses it later.
Panel C (bottom): Iran vs. China—scale and regime change.
Iran and China start the 1960s–1970s in a broadly comparable order of magnitude; after 1979 their paths bifurcate. Iran experiences a trough and slow rebuild, while China undergoes a two-stage regime shift: a steepening in the late-1980s/1990s and an even steeper rise post-2000 that lifts the level by several orders of magnitude on the log scale. The persistent slope gap through the 1990s–2010s makes absolute catch-up arithmetically implausible for Iran even with robust 2000s growth.
Synthesis.
Across panels, Figure 10 shows that (i) Iran’s pre-1979 slope was competitive, (ii) the post-1979 regime shift raised its implied doubling time for roughly a decade and a half, and (iii) peers that maintained low doubling times (the Tigers; the Netherlands/Israel; later China) compounded into permanently higher levels. The resulting divergence is thus a consequence of rate differentials sustained over long horizons, not merely one-off level shocks.
3.7 Per-capita capacity (Figure 11)
Normalizing by population clarifies that the Asian Tigers exhibit early inflection and multi-decadal compounding on a per-capita basis, reflecting institutional deepening (graduate training, research careers, internationalization). Iran’s post-1979 trough is stark on this metric: per-capita output falls sharply, then climbs in the 2000s, but the relative gap with Tigers and mature systems (Netherlands, Israel) remains large. The normalization shows Iran’s 2000s recovery was driven primarily by scale rebuilding rather than a per-capita surge sufficient for convergence.
3.8 Scientific growth momentum: implied years to double output (Figure 12)
Figure 12 tracks implied years to double publication output—a momentum metric computed from rolling growth rates (lower values indicate stronger compounding). The vertical dashed line marks 1979.
Figure 12: Growth momentum and implied doubling times
Panel A (top): Iran vs. the Asian Tigers.
Iran enters the 1970s with comparatively low doubling times (about 3–5 years), signalling a live, fast-growing system. Around 1979–1981 the red curve spikes abruptly to well above 20 years, a signature of collapse/stagnation; this regime shift coincides with the revolution and its immediate sequelae. Momentum then gradually normalizes in the late 1980s (back toward $\sim$6–7 years), briefly improves around 2000 (approaching $\sim$3 years), and subsequently deteriorates again during the 2010s (rising toward $\sim$8–10 years).
In sharp contrast, South Korea and Taiwan maintain persistently low doubling times through the 1980s–1990s (roughly $\sim$3–6 years), exactly the regime required for multi-decadal exponential scale-up. Singapore shows a similar pattern: momentum strengthens through the 1980s–1990s, then lengthens in the 2010s as a maturing system naturally slows. The key message of Panel A is that the Tigers sustain low doubling times for long horizons just as Iran transitions into a high-doubling-time regime circa 1979; even Iran’s later recovery does not keep momentum low for long enough to deliver convergence.
@@ 3.8 Scientific growth momentum: implied years to double output (Figure 12)
The Netherlands and Israel display the characteristic profile of mature research ecosystems: doubling times hover in the low-teens and gradually lengthen over time, reflecting steady compounding from a high base rather than explosive growth from a low base. Greece climbs from mid-single-digit doubling times in the 1960s toward the teens by the 1990s and 2000s, consistent with long-horizon capacity building in a smaller system. Against this backdrop, Iran’s trajectory is distinctive: a pre-1979 low-doubling-time regime (fast momentum), a dramatic spike to >>20 years at the turning point, a late-1980s normalization, a short early-200
@@ 3.9 Counterfactual Analysis: Quantifying the Knowledge Deficit
To quantify the magnitude of the divergence, we developed a suite of counterfactual models, each designed to project Iran’s scientific output under a scenario of uninterrupted growth. These models range from simple mathematical extrapolations to more sophisticated, empirically grounded simulations. The results, visualized in Figure 13, collectively illustrate the immense opportunity cost of the post-revolutionary disruption.
@@ 3.9.1 Model B (Perpetual Exponential Growth)
This model serves as a theoretical benchmark to illustrate the raw power of compounding. It assumes that the institutional momentum Iran had built up in the decade prior to the revolution would continue indefinitely. We calculate the Compound Annual Growth Rate (CAGR) from 1970 to 1978, which stood at a robust $r\approx 13.1\%$. The projection is then calculated using the standard exponential growth formula:
@@
As shown by the dotted gray line in Figure 13, this model projects an unrealistic output of over 1.2 million papers by 2024. While no system sustains such a high growth rate forever, it provides a startling mathematical ceiling for Iran’s lost potential, suggesting a trajectory that could have theoretically rivaled the absolute scale of modern-day China.
@@ 3.9.2 Model C (Logistic S-Curve Growth with Carrying Capacity)
This model introduces a crucial element of realism: the concept of maturation. Scientific systems, like all growth processes, eventually face diminishing returns as they exhaust available resources (e.g., funding, human capital, institutional capacity), causing their growth to slow and follow an S-shaped curve. The logistic model captures this by incorporating a ”carrying capacity” ($K$), which represents the plausible upper limit of a nation’s scientific system.
Justification for Carrying Capacity (K): We set $K=150,000$ publications per year. This figure is not arbitrary; it is chosen to represent the scale of a mature, top-tier, and stable scientific nation, comparable to the current output of countries like South Korea, Italy, or Canada. In the context of scientometrics, K is a function of sustained investment, population size, and effective science policy. Nations achieving this scale of output typically maintain a Gross Expenditure on R&D (GERD) consistently above 2% of GDP, coupled with a high density of researchers within the population. The value of 150,000 thus represents a plausible ceiling that a nation with Iran’s demographic potential and pre-1979 momentum could have realistically achieved, had it maintained stability and adopted similar long-term investment strategies.
@@
The dashed blue line in Figure 13 shows this path. Iran’s growth is initially rapid, but then gracefully slows as it approaches the 150,000-publication mark. This model projects a far more plausible scenario where Iran would have become a major, stable scientific power.
@@ 3.9.3 Model D (South Korea as a Growth-Proxy)
This final model is arguably the most compelling as it is grounded in a real-world success story. It answers the question: ”What if Iran, starting from its superior position in 1978, had simply followed South Korea’s subsequent path of development?”
@@
The result, shown by the dashed green line in Figure 13, is remarkable. It suggests that Iran would not only have matched but significantly surpassed South Korea, reaching an annual output of approximately **245,000 papers** by 2024. This is because Iran would have been applying Korea’s high growth rates to its own, larger initial base of scientific output. This data-driven simulation provides a powerful, tangible estimate of the ”knowledge deficit.”
@@
Collectively, the vast chasm between the solid red line (Iran’s actual path) and the plausible scenarios modeled by the blue and green lines provides a robust, data-grounded estimate of the profound and lasting cost of the 1979 disruption.
4 Discussion and Conclusion
The empirical evidence reveals a persistent and multifaceted divergence in Iran’s scientific trajectory following the 1979 Revolution. The ”lost decade” was not merely a pause in publication growth; it was a critical period where the compounding growth of the Asian Tigers—in both quantity and quality—created an unbridgeable gap. Our analysis shows that the disruption’s primary cost was the loss of compounding momentum. While Iran’s scientific community has shown remarkable resilience in rebuilding its output, it has been fighting a battle against arithmetic, trying to close a gap with nations that benefited from decades of uninterrupted growth fueled by stability and strategic investment.
The story, however, is more complex than volume alone. The data on FWCI and top-percentile papers suggests that the “knowledge deficit” quantified by our counterfactual models extends beyond publication counts to a deficit in scientific influence. Iran’s recovery has been characterized by achieving global scale but not yet global leadership in impact. While a lag between quantity and quality is common in rapidly expanding systems, it is particularly pronounced here and likely stems from a confluence of structural factors.
This persistent quality gap is likely driven by mechanisms operating at both the international and domestic levels. Internationally, decades of relative isolation, exacerbated by geopolitical constraints and sanctions, have significantly constrained opportunities for deep collaboration with leading global research groups. As international collaboration is a key driver of FWCI, reduced integration into elite networks inherently limits knowledge transfer and citation visibility. Domestically, science policy and university promotion criteria during the rapid recovery phase often prioritized quantitative metrics over impact. This incentive structure may inadvertently favor volume maximization strategies—such as publishing in lower-tier venues or engaging in incremental research—at the expense of the long-term, resource-intensive projects that typically yield high-impact breakthroughs. Consequently, the deficit in scientific influence reflects not only lost time but the enduring structural challenges of re-integrating a large-scale system into the highest echelons of global science.
The key lesson from this comparative analysis is the critical role of long-term stability, sustained investment, and international integration in building a national scientific enterprise. Short-term disruptions have multi-generational consequences, affecting not just the number of papers produced but the depth of their integration into the global conversation of science. The story of Iran and the Asian Tigers is a stark illustration of this fundamental principle: the dividends of stability are immense, and the costs of disruption are generational, echoing through both the size and the influence of a nation’s scientific voice. These findings highlight that sustained investment in scientific institutions and global collaboration networks is essential for resilience after political disruption.

### v2

@@ 3.6 Cross-country synthesis on a log scale
Across all panels, the data demonstrates that (i) Iran’s pre-1979 slope was competitive, (ii) the post-1979 regime shift dramatically flattened this slope, raising its implied doubling time for roughly a decade and a half, and (iii) peers that maintained steep, low-doubling-time regimes (the Tigers; the Netherlands/Israel; later China) compounded into permanently higher levels. The resulting divergence is thus a consequence of rate differentials sustained over long horizons, not merely one-off level shocks.
This divergence in compounding growth is evident across all peer groups. In the comparison with the ”Asian Tigers” (Panel A), Iran’s pre-1979 momentum is clear; its slope is competitive—at times steeper than South Korea’s—and its level is comparable to Taiwan’s. Immediately after 1979, Iran’s curve flattens and dips, signalling a collapse in compounding. By contrast, South Korea, Taiwan, and Singapore enter multi-decadal high-slope regimes during the 1980s–1990s. This fundamental divergence in rates explains why Iran’s strong 2000s recovery does not translate into convergence.
Similarly, when compared to established systems (Panel B), the Netherlands and Israel trace mature, high-level trajectories with moderately rising slopes, consistent with large, stable research systems. Relative to this group, Iran’s post-1979 dip is distinctive: the level falls below Greece in the 1980s and only re-approaches the regional band after 2000. The key contrast is regime stability: the mature systems preserve low implied doubling times throughout the 1980s–1990s, whereas Iran shifts into a high-doubling-time regime.
Finally, the comparison with China (Panel C) illustrates a bifurcation of paths. While Iran experiences a trough and slow rebuild, China undergoes a two-stage regime shift: a steepening in the 1990s and an even steeper rise post-2000. The persistent slope gap through the 1990s–2010s makes absolute catch-up arithmetically implausible for Iran, even with its robust 2000s growth.
3.7 Per-capita capacity
Normalizing by population (Figure 12) clarifies that the Asian Tigers exhibit early inflection and multi-decadal compounding on a per-capita basis, reflecting institutional deepening (graduate training, research careers, internationalization). Iran’s post-1979 trough is stark on this metric: per-capita output falls sharply, then climbs in the 2000s, but the relative gap with Tigers and mature systems (Netherlands, Israel) remains large. The normalization shows Iran’s 2000s recovery was driven primarily by scale rebuilding rather than a per-capita surge sufficient for convergence.
3.8 Scientific growth momentum: implied years to double output
The Figure tracks implied years to double publication output—a momentum metric computed from rolling growth rates (lower values indicate stronger compounding). The vertical dashed line marks 1979.
Figure 13: Growth momentum and implied doubling times.
@@ 3.8 Scientific growth momentum: implied years to double output
Iran enters the 1970s with comparatively low doubling times (about 3–5 years), signalling an active, fast-growing system. Around 1979–1981 the red curve spikes abruptly to well above 20 years, a signature of stagnation; this regime shift coincides with the revolution and its immediate sequelae. Momentum then gradually normalizes in the late 1980s (back toward $\sim$6–7 years), briefly improves around 2000 (approaching $\sim$3 years), and subsequently deteriorates again during the 2010s (rising toward $\sim$8–10 years).
In sharp contrast, South Korea and Taiwan maintain persistently low doubling times through the 1980s–1990s (roughly $\sim$3–6 years), exactly the regime required for multi-decadal exponential scale-up. Singapore shows a similar pattern: momentum strengthens through the 1980s–1990s, then lengthens in the 2010s as a maturing system naturally slows. Panel A shows that the Tigers sustain low doubling times for long horizons just as Iran transitions into a high-doubling-time regime circa 1979; even Iran’s later recovery does not keep momentum low for long enough to deliver convergence.
@@
The Netherlands and Israel display the characteristic profile of mature research ecosystems: doubling times hover in the low-teens and gradually lengthen over time, reflecting steady compounding from a high base rather than explosive growth from a low base. Greece climbs from mid-single-digit doubling times in the 1960s toward the teens by the 1990s and 2000s, consistent with long-horizon capacity building in a smaller system. Against this backdrop, Iran’s trajectory is distinctive: a pre-1979 low-doubling-time regime (fast momentum), a dramatic spike to >>20 years at the turning point, a late-1980s normalization, a short early-2000s before slowing again.
@@ 3.9 Counterfactual Analysis: Quantifying the Knowledge Deficit
To quantify the magnitude of the divergence, we developed a suite of counterfactual models, each designed to project Iran’s scientific output under a scenario of uninterrupted growth. These models range from simple mathematical extrapolations to more sophisticated, empirically grounded simulations. The results, visualized in Figure 14, collectively illustrate the significant opportunity cost of the post-revolutionary disruption.
@@ 3.9.1 Model B (Perpetual Exponential Growth)
This model serves as a theoretical benchmark to illustrate the power of compounding. It assumes that the institutional momentum Iran had built up in the decade prior to the revolution would continue indefinitely. We calculate the Compound Annual Growth Rate (CAGR) from 1970 to 1978, which stood at a robust $r\approx 13.1\%$. The projection is then calculated using the standard exponential growth formula:
@@
As shown by the dotted gray line in Figure 14, this model projects an unrealistic output of over 1.2 million papers by 2024. While no system can sustain such a high growth rate forever, it highlights a significant mathematical limitation for Iran’s lost potential, suggesting a trajectory that could have theoretically rivaled the scale of modern-day China.
@@ 3.9.2 Model C (Logistic S-Curve Growth with Carrying Capacity)
This model introduces a significant element of realism: the concept of maturation. Scientific systems, like all growth processes, eventually face diminishing returns as they exhaust available resources (e.g., funding, human capital, institutional capacity), causing their growth to slow and follow an S-shaped curve. The logistic model captures this by incorporating a ”carrying capacity” ($K$), which represents the plausible upper limit of a nation’s scientific system.
Justification for Carrying Capacity (K): We set $K=150,000$ publications per year. This figure is not arbitrary; it is chosen to represent the scale of a mature, top-tier, and stable scientific nation, comparable to the current output of countries like South Korea, Italy, or Canada. In the context of scientometrics, K is a function of sustained investment, population size, and effective science policy. Nations achieving this scale of output typically maintain a Gross Expenditure on R&D (GERD) consistently above 2% of GDP, coupled with a high density of researchers within the population. The value of 150,000 thus represents a plausible limitation that a nation with Iran’s demographic potential and pre-1979 momentum could have realistically achieved, had it maintained stability and adopted similar long-term investment strategies.
@@
The dashed blue line in Figure 14 shows this path. Iran’s growth is initially rapid, but then subsequently slows as it approaches the 150,000-publication mark. This model projects a far more plausible scenario where Iran would have become a major, stable scientific power.
@@ 3.9.3 Model D (South Korea as a Growth-Proxy)
This final model is compelling as it is grounded in a real-world success story. It answers the question: ”What if Iran, starting from its superior position in 1978, had simply followed South Korea’s subsequent path of development?”
@@
The result, as shown by the dashed green line in Figure 14, indicates that Iran would not only match but significantly surpass South Korea, reaching an annual output of approximately 245,000 papers by 2024. This is because Iran would have been applying Korea’s high growth rates to its own, larger initial base of scientific output. This data-driven simulation provides a significant, tangible estimate of the ”knowledge deficit.”
@@
Collectively, the large gap between the solid red line (Iran’s actual path) and the plausible scenarios modeled by the blue and green lines provides a robust, data-grounded estimate of the significant and lasting cost of the 1979 disruption.
3.10 Comparing Iran to Two Empirical Counterfactuals
Figure 15 extends the counterfactual analysis introduced in Figure 14 by incorporating two empirically grounded benchmarks: (i) a Synthetic Iran constructed via the Synthetic Control Method (SCM), and (ii) a scaled version of South Korea’s historical trajectory calibrated to Iran’s pre-1979 level. Together, these models allow us to quantify both a conservative and an ambitious estimate of Iran’s “knowledge deficit” by 2024.
(a) Scientific Output: Actual Iran vs. Synthetic and Scaled-Korea Iran.
(b) Gap in scientific output relative to Synthetic Iran (SCM) and scaled-Korea counterfactuals.
(c) Cumulative scientific output under actual and counterfactual trajectories.
Figure 15: Comparative evaluation of Iran’s scientific output under two data-driven counterfactuals.
It is important to clarify how the modelling strategy in Figure 14 differs from the synthetic–control framework adopted in Figure 15. Figure 14 presents three data–driven but explicitly model–based counterfactual trajectories. All three paths rely on real empirical information, but they remain stylized projections designed to illustrate a range of plausible, data–anchored futures.
By contrast, Figure 15 implements a fully non–parametric, comparison–based approach using the Synthetic Control Method (SCM). Rather than assuming a functional form or imposing a specific growth structure, SCM constructs a counterfactual Iran directly from the weighted combination of real countries that most closely match Iran’s pre–1979 economic and educational profile. Whereas the models in Figure 14 illustrate how Iran might have evolved under different empirically informed assumptions, the SCM in Figure 15 provides a data–driven comparative benchmark grounded solely in observed pre–treatment similarities. Together, these complementary approaches allow us to triangulate the magnitude of Iran’s post–1979 divergence using both model–based and comparison–based counterfactual evidence.
The SCM is a methodology widely used in economics and political science for estimating the causal impact of major historical shocks [abadie2010synthetic]. The core idea of SCM is to approximate the “treated unit”—here, Iran after 1979—with a weighted combination of untreated comparison countries whose pre–intervention characteristics closely resemble those of the treated unit. Unlike traditional regression methods, SCM does not impose a functional form; instead, it builds the closest possible replica of Iran’s pre–1979 trajectory using convex weights that are optimized to minimize the pre–treatment prediction error.
In this study, the SCM model is constructed using annual publication counts from Scopus as the dependent variable and four structural predictors from the World Bank database: GDP per capita (constant prices), tertiary–education enrollment, government expenditure, and trade openness. The model is trained strictly on the pre–1979 period, ensuring that the synthetic version of Iran is calibrated exclusively to the historical conditions that existed before the Revolution.
The optimization procedure assigns a weight of 1.0 to South Korea as the sole donor country. This outcome is consistent with the data structure of our predictor set: among all candidate countries, Korea is the only one whose pre–1979 trajectory in per–capita GDP, tertiary enrollment, public expenditure, and trade openness jointly reproduces Iran’s pre–revolution profile with minimal prediction error. In other words, the SCM algorithm selects Korea not because of any arbitrary modeling choice, but because it is the unique country in the donor pool whose multidimensional economic–education vector most closely approximates Iran’s historical values. This makes Korea the statistically optimal donor unit for constructing a credible counterfactual.
The resulting “Synthetic Iran” represents an empirical estimate of how Iran’s scientific output would likely have evolved in the absence of the 1979 disruption, assuming its pre–Revolution structural fundamentals had remained stable. This counterfactual serves as a conservative benchmark for evaluating the long–term impact of the Revolution and the prolonged interruption of higher–education institutions on Iran’s scientific productivity.
Panel (a) shows that prior to 1979, all three trajectories follow a similar rising path, reflecting the shared developmental momentum of Iran and the Asian Tigers. Immediately after the Revolution and the Cultural Revolution–era shutdown of universities, the observed curve collapses while both counterfactual trajectories continue rising smoothly. The SCM path—which uses only economic and educational predictors in the pre-1979 period—grows moderately and represents a lower-bound estimate of Iran’s potential. The scaled-Korea curve grows substantially faster, reflecting the historically high compounding rates achieved by South Korea during 1980-2020.
Panel (b) quantifies the annual “gap” between actual output and each counterfactual. Both gaps remain close to zero before 1979 but widen significantly afterward. By the early 2000s the deficit becomes structurally embedded, and by 2024 the gap relative to the SCM counterfactual reaches roughly 22,847 articles, while the gap relative to the scaled-Korea benchmark exceeds 107,889. This clearly demonstrates that Iran’s post-1979 recovery never closed the structural distance accumulated during the lost decade.
Panel (c) displays cumulative output—a metric that captures the compounding nature of scientific production capacity. Whereas Iran’s cumulative total reaches approximately 864,675 papers by 2024, the SCM counterfactual produces around 1.42 million, and the scaled-Korea model exceeds 2.71 million. The cumulative “knowledge deficit” therefore ranges from a conservative estimate of 551,129 lost publications (SCM) to a substantially larger deficit of 1.85 million (scaled-Korea). The large divergence in cumulative terms underscores that the primary long-term cost of the 1979 disruption was not solely the immediate collapse of output, but the permanent loss of compounding scientific momentum.
In sum, Figure 15 demonstrates that even under the most conservative assumptions, Iran’s scientific system today represents only a fraction of its plausible counterfactual potential. The SCM model anchors a lower-bound scenario based on structural predictors, while the scaled-Korea trajectory approximates a realistic upper-bound scenario consistent with the growth path of Iran’s closest developmental analogue. Both converge on the same conclusion: the Revolution-induced shock resulted in a generational loss of Iran’s scientific capacity.
4 Conclusion
This paper provides a comprehensive, data–driven reconstruction of Iran’s scientific trajectory over more than six decades and evaluates its long–term evolution through two independent counterfactual frameworks. The findings demonstrate that the disruption triggered by the 1979 Revolution produced a structural break whose effects extend beyond the immediate decline in scientific activity. Rather than representing a temporary interruption, the post–1979 decade reset the underlying growth rate of Iran’s scientific system and altered its long–run development path.
The counterfactual analyses clarify the magnitude of this divergence. The Synthetic Control Method, which selects South Korea as the single optimal pre–1979 match for Iran, indicates that Iran’s scientific output would have followed a substantially steeper growth trajectory in the absence of institutional upheaval. A second model, constructed by scaling Korea’s post–1979 trajectory to Iran’s 1979 baseline, yields an even more accelerated path. Both counterfactuals converge on the same conclusion: by 2024, Iran’s actual scientific output remains far below the level that would have been expected had its pre–1979 momentum continued uninterrupted.
The comparison between these reconstructed trajectories and Iran’s observed performance highlights a key insight: Iran’s scientific system successfully regained scale in the post–2000 period, but it did so along a shallower curve than its historical peers. The resulting “knowledge deficit” is therefore not merely the product of lost years, but the cumulative consequence of a lost growth rate. The analysis shows that once a nation’s scientific institutions lose continuity, the long–term effects compound—even in the presence of later recovery efforts.
In sum, the evidence demonstrates that Iran’s scientific development has been shaped not only by its capacity to produce publications but by the structural conditions that enable sustained, compounding growth. The counterfactual estimates presented in this study quantify the generational cost of the disruption that occurred in 1979 and underscore the central role of stability, openness, and institutional continuity in shaping national scientific trajectories. This suggests that reversing the long-term deficit identified here is contingent upon rebuilding these foundational conditions and re-establishing a growth environment capable of restoring the momentum lost more than four decades ago.
Data Availability
The datasets generated and/or analyzed during the current study are available from the corresponding author on reasonable request.